# Spectro-temporal analysis of ultra-fast radio bursts using per-channel arrival times

### v2

@@ Abstract
Fast radio bursts (FRBs), especially those from repeating sources, exhibit a rich variety of morphologies in their dynamic spectra (or waterfalls). Characterizing these morphologies and spectro-temporal properties is a key strategy in investigating the underlying unknown emission mechanism of FRBs. This type of analysis has been typically accomplished using two-dimensional Gaussian techniques and the autocorrelation function (ACF) of the waterfall. These techniques are effective and precise at all duration scales, but can be limited in the presence of scattered tails, complex morphologies, or recently observed microshot forests. Here, we present a technique that involves the tagging of per-channel arrival times of an FRB to perform spectro-temporal measurements using a Gaussian profile model for each channel. While scattering and dispersion remain important and often dominating sources of uncertainty in measurements, this technique provides an adaptable and firm foundation for obtaining spectro-temporal properties from all types of FRB morphologies. We present measurements using this technique of several hundred bursts across 12 repeating sources, including over 400 bursts from the repeating sources FRB 20121102A, FRB 20220912A, and FRB 20200120E, all of which exhibit recently observed microsecond-long ultra-FRBs, as well as 143 drift rates. In addition to retrieving the known relationship between sub-burst slope and duration, we explore other correlations between burst properties. We find that the sub-burst slope law extends smoothly to ultra-FRBs, and that ultra-FRBs appear to form a distinct population in the duration-frequency relation.
@@ 2.1 Arrival times pipeline
An additional step is performed in the presence of multiple components to measure the drift rate between components. Using the center frequency and the time of each component, obtained from the integrated 1D spectrum and time series, a line analogous to that in Equation 3 is fitted to define the drift rate . We measure the ‘total duration’ of an FRB with multiple components, which may have unpredictable spacings between them, as the time difference between the first and last components. This estimate of the duration turns out to be sufficient for our purposes, though other definitions may be required depending on the research goals.
@@ 3 Observations and Data
The data used for this study was obtained from previously published observational studies. The majority of the bursts analysed here have not had measurements of their sub-burst slopes performed before, and we also reanalyse the spectro-temporal properties of many bursts from earlier studies. To the best of our knowledge, the ultra-FRBs analysed here represent the first sub-burst slope measurements for these extremely short bursts. This section provides a brief summary of the repeating FRBs included in this study, with a particular focus on those exhibiting ultra-FRBs or bursts with microsecond-long durations (specifically, FRB 20121102A, FRB 20220912A, and FRB 20200120E). Additionally, we outline the properties of the bursts analysed and describe the DM strategy employed during the measurements. In general, we try to use the DM of the shortest burst in a cohort and apply it to all the bursts when performing measurements. Later, we recalculate the DM of each burst individually and repeat our measurements to assess any major differences. As in earlier studies (Chamma et al., 2023; Brown et al., 2024), we process and convert all data from various formats (e.g., PSRCHIVE, PSRFITS, filterbank, etc.) into Python numpy (.npz) format for analysis. The details of this conversion are described here, with reference to previous works where appropriate.
@@ 3.1.1 FRB 20121102A
For this dataset we investigated applying the DM of one of the shortest bursts from the source (in this case 560.105 pc/cm3 from the ultra-FRB B30 of Snelders et al. 2023) to the majority of the data. This approach was chosen because the DM of the shortest burst is likely to have high precision due to its sharp profile and high frequency. Jahns et al. (2023) also recommend this step when analysing bursts emitted within weeks of each other, as their results suggest that short-term DM variations are caused by intra-burst drift. The data from Hewitt et al. (2022) were not included in this step, and these data were analysed with their applied DM of 560.5 pc/cm3. We hypothesized that the DM of the shortest burst could be a highly precise measurement of a source’s DM, which would be indicated by an improved agreement with fits to the sub-burst slope law. We did indeed observe slight improvements in the goodness of fit compared to using each burst’s individual DM. A more thorough investigation into optimizing the DM would also need to account for the source’s DM evolution over time, which we decided to be beyond the scope of this paper. The results of these DM investigations are discussed further in Appendix C.
@@ 3.2 Burst DM Measurements
As mentioned, to assess any major differences in our results due to the choice of DM, we repeat our measurements at the individually determined DM of each burst. We repeated all spectro-temporal measurements at the burst DM reported by the original authors, if there was one. For bursts without an individually calculated DM, we used DM_phase555https://github.com/danielemichilli/DM_phase (CHIME/FRB Collaboration et al., 2019; Seymour et al., 2019) to recalculate the DM using an algorithm that preserves structure by maximizing coherent power instead of S/N, and repeated the measurement at that DM. The datasets whose burst DMs we recomputed were those from Gajjar et al. (2018); Michilli et al. (2018); Oostrum et al. (2020); Scholz et al. (2016); Snelders et al. (2023); Hewitt et al. (2022); Zhang et al. (2023); Nimmo et al. (2022, 2023); Pleunis et al. (2021b); Marthi et al. (2020); Bethapudi et al. (2023). For most of the FRBs we recomputed DMs for, we searched for optimal DMs in a range of 40 pc/cm3 from the applied DM in steps of 0.1 pc/cm3. For shorter FRBs (namely those from Snelders et al. 2023 and Nimmo et al. 2022, 2023, as well as bursts from FRB 20180916B), a finer grid was needed over a smaller range, so we searched DMs 10 pc/cm3 from the applied DM in steps of 0.01 pc/cm3. For the remaining datasets that reported individual burst DMs, we repeated our measurements at those DMs.
@@ 4.1 Spectro-temporal relationships
Table 2 presents the results of these fits for each FRB source analysed. The fit for FRB 20121102A, which includes 149 bursts measured and covers the widest range of frequencies and durations, yields and ms. The uncertainties here and throughout (including figures) are quoted at the 1 level. This fit is shown as the black dash-dot line in Figure 4. Note that the curved appearance of the fit line results from the logarithmic scale used on the duration axis. Due to the general form of the fit, agreement with Equation 4 is qualitatively captured by the magnitude of (as well as ), where a value near zero indicates good agreement and larger values suggest otherwise. Assuming that must be zero (i.e., that the sub-burst relation holds true) allows to serve as an indicator of data quality. A large may result from a limited variety or incomplete sample of bursts in terms of frequency and/or duration. This trend is evident in Table 2 for the remaining sources, where fluctuates around the value found for FRB 20121102A, while generally increases significantly as the number of available bursts decreases. FRB 20220912A challenges this trend since is large despite the large number of bursts, however this may be due to an actual deviation from the sub-burst slope law for this source.
@@ 4.1.2 Relations between , , and
The – plot appears to exhibit a linear correlation, as was found with previous results obtained over broad frequency ranges but with fewer bursts (Houde et al., 2019; Chamma et al., 2023; Brown et al., 2024); however, there again remains significant statistical scatter. We fit a line to the complete set of measurements and obtain . Previous values for this correlation are around , after inverting and scaling from the FWHM for comparison here. Despite the present value being nearly double, we note that, because of the spread in measurements, either fit could be visually attributed to the data.
Two fits are overlaid on the – plot of the form . The blue and orange lines are fit to all bursts with durations longer and shorter than 300 s, respectively. We find msMHz for the longer FRBs, and msMHz for the ultra-FRBs. Despite the statistical scatter in measurements, the blue line reflects well the trend of measurements, especially the very low frequency bursts from FRB 20180916B. The fit to the ultra-FRBs (orange) also describes the shortest of bursts well. Bursts at around 1.4 GHz, especially from the sources FRB 20200120E and FRB 20220912A, connect the duration gap between the two fits, suggesting a continuum may exist. In that case, the blue and orange fits may represent opposite ends of an envelope. The functional dependence is the prediction of the TRDM and consistent with results reported in Chamma et al. (2023). There is room for other interpretations and functional forms that may describe this data (such as ). However, we focus here on the context that the TRDM provides and follow predictions made by Kumar et al. (2024), which investigates modifications to the TRDM at short durations and under various propagation effects. In Kumar et al. (2024), bursts that obey the sub-burst slope law and vary in duration by several orders of magnitude can show the sort of separation reflected in the two fits shown.
@@ 4.1.3 Changes when measuring at individual burst DMs
Generally, we observe an expected shift of slope measurements upwards towards positive slopes. After applying the same measurement filters applied to the earlier measurements (Section 2.2), 325 sub-burst measurements remain. The reduction in burst count is primarily due to the increase in uncertainty for values of d/d that are close to or overlap with 0 ms/MHz. Since the burst DM is found using an algorithm that generally aligns bursts vertically, this situation is more common and we observe a corresponding decrease in usable measurements. Nonetheless, we still observe a downwards trend with increasing duration consistent with the sub-burst slope relation. The fit parameters obtained for the sub-burst slope relation for bursts from FRB 20121102A was and ms, as well as a larger reduced-. For the remaining sources, the fit parameters found showed similar differences, with the reduced- either similar or much larger compared to those tabulated in Table 2.
@@ 4.2 Drift rates
Figure 6 shows the drift rate measurements, each multiplied by the mean frequency of their underlying components, plotted against their duration. Overlaid is the fit obtained for the analogous plot of the sub-burst slopes in Figure 4. We observe good agreement between this fit and the drift rate trend, supporting the findings in Jahns et al. (2023) and Brown et al. (2024) that these two measurements follow similar trends. Interestingly, the positive drifts approximately follow the negative of this relationship, shown with the faded line, though the small number of points makes it difficult to conclude for certain. This suggests a connection with the negative drifts and the existence of a sign change (Rajabi et al., 2020), which can explain this aspect of the positive drifts. To double check this result, we repeated all our drift rate measurements using the ACF method. We obtained 132 drift rates from our sample in this way. Using the same durations, we saw a nearly identical trend in the measurements, despite some differences in measured drifts.
@@ 5 Discussion
The slope–frequency relation is a strong correlation that is observed in our results here and predicted in the TRDM (Eq. 7 of Chamma et al., 2023). Despite three outliers, the addition of low frequency LOFAR data for FRB 20180916B seems to suggest a shared trend between at least that source and FRB 20121102A. This is because the same scaling of the fit shown in the – panel of Figure 5 describes the FRB 20180916B data and indeed the data from all sources well. In the TRDM the scaling constant for this relationship is equivalent to where is the rest frame frequency of emission of the FRB. This relationship is also expected in the context of curvature radiation by charged bunches in the magnetosphere of magnetars (Wang et al., 2022). In this vein, a recent scintillation analysis for FRB 20221022A argues that the lateral size of the FRB emission region in that source to be less than km (Nimmo et al., 2025).
@@
We also note that the sub-burst slope of ultra-FRBs is especially sensitive to dispersion errors. Because of their extremely short duration, ultra-FRBs appear close to vertical on a perfectly dedispersed waterfall. It follows that even small levels of over-dedispersion can change the sign of their sub-burst slope (Kumar et al., 2024, their fig. 10). In principle, this could explain the tendency for some of the sub-burst slopes of shorter bursts, especially those from FRB 20200120E (durations between and ms) and FRB 20121102A and FRB 20220912A (durations above ms), to ‘bump’ upwards in Figure 4.
@@ 6 Summary and Conclusions
We have developed and applied a measurement pipeline that uses the arrival times of an FRB pulse in each frequency channel of its dynamic spectrum to perform sub-burst slope/intra-burst drift measurements. The pipeline successfully delivered high precision measurements of the slope including for ultra-FRBs. In most cases, earlier 2D Gaussian techniques yielded measurements with similar values and uncertainties, whereas the ACF method yielded the smallest uncertainties across all methods used. Drift rates were also measured for bursts with multiple components, yielding a total of 143 measurements.
@@ Acknowledgements
The authors would like to thank the anonymous referee, whose feedback greatly improved the quality of the manuscript and its conclusions.
@@ Appendix C DM optimization using the sub-burst slope relation
We apply this process for the sources FRB 20121102A and FRB 20220912A separately. Note that we exclude the many bursts from Hewitt et al. (2022) for FRB 20121102A due to their different applied starting DM and burst B1 from Hewitt et al. (2023) for FRB20220912A due to the presence of significant blending. The starting DM used is the DM of the shortest (or nearly the shortest) burst from that source. In the case of FRB 20121102A that burst is the microburst B30 of Snelders et al. (2023), and burst B1 of Hewitt et al. (2023) for FRB 20220912A. The results of this are shown in Table 4.

### v3

@@ Abstract
Fast radio bursts (FRBs), especially those from repeating sources, exhibit a rich variety of morphologies in their dynamic spectra (or waterfalls). Characterizing these morphologies and spectro-temporal properties is a key strategy in investigating the underlying unknown emission mechanism of FRBs. This type of analysis has been typically accomplished using two-dimensional Gaussian techniques and the autocorrelation function (ACF) of the waterfall. These techniques are effective and precise at all duration scales, but can be limited in the presence of scattered tails, complex morphologies, or recently observed microshot forests. Here, we present a technique that involves the tagging of per-channel arrival times of an FRB to perform spectro-temporal measurements using a Gaussian profile model for each channel. While scattering and dispersion remain important and often dominating sources of uncertainty in measurements, this technique provides an adaptable and firm foundation for obtaining spectro-temporal properties from all types of FRB morphologies. We present measurements using this technique of several hundred bursts across 12 repeating sources, including over 400 bursts from the repeating sources FRB 20121102A, FRB 20220912A, and FRB 20200120E, all of which exhibit recently observed microsecond-long ultra-FRBs, as well as 143 multi-component drift rates. In addition to retrieving the known relationship between sub-burst slope and duration, we explore other correlations between burst properties. We find that the sub-burst slope law extends smoothly to ultra-FRBs, and that ultra-FRBs appear to form a distinct population in the duration-frequency relation.
@@ 2.1 Arrival times pipeline
The burst duration measured from the integrated time-series can be artificially broadened by drifting. We can alternatively define the burst duration as the average of the ’s found in each frequency channel with sufficient S/N (as defined in step 5). However, this may be limited by the S/N and the morphology of the burst. We also discuss using the measured slope to ‘correct’ the duration by removing residual drift in Appendix C.
An additional step is performed in the presence of multiple burst components to measure the drift rate between components. Using the center frequency and the time of each component, obtained from the integrated 1D spectrum and time series, a line analogous to that in Equation 3 is fitted to define the drift rate . We measure the ‘total duration’ of an FRB with multiple components, which may have unpredictable spacings between them, as the time difference between the first and last components. This estimate of the duration turns out to be sufficient for our purposes, though other definitions may be required depending on the research goals.
@@ 3 Observations and Data
The data used for this study was obtained from previously published observational studies. The majority of the bursts analysed here have not had measurements of their sub-burst slopes performed before, and we also reanalyse the spectro-temporal properties of many bursts from earlier studies. To the best of our knowledge, the ultra-FRBs analysed here represent the first sub-burst slope measurements for these extremely short duration bursts. This section provides a brief summary of the repeating FRBs included in this study, with a particular focus on those exhibiting ultra-FRBs or bursts with microsecond-long durations (specifically, FRB 20121102A, FRB 20220912A, and FRB 20200120E). Additionally, we outline the properties of the bursts analysed and describe the DM strategy employed during the measurements. In general, we try to use the DM of the shortest duration burst in a cohort and apply it to all the bursts when performing measurements. Later, we recalculate the DM of each burst individually and repeat our measurements to assess any major differences. As in earlier studies (Chamma et al., 2023; Brown et al., 2024), we process and convert all data from various formats (e.g., PSRCHIVE, PSRFITS, filterbank, etc.) into Python numpy (.npz) format for analysis. The details of this conversion are described here, with reference to previous works where appropriate.
@@ 3.1.1 FRB 20121102A
For this dataset we investigated applying the DM of one of the shortest bursts from the source (in this case 560.105 pc/cm3 from the ultra-FRB B30 of Snelders et al. 2023) to the majority of the data. This approach was chosen because the DM of the shortest duration burst is likely to have high precision due to its sharp profile and high frequency. Jahns et al. (2023) also recommend this step when analysing bursts emitted within weeks of each other, as their results suggest that short-term DM variations are caused by intra-burst drift. The data from Hewitt et al. (2022) were not included in this step, and these data were analysed with their applied DM of 560.5 pc/cm3. We hypothesized that the DM of the shortest duration burst could be a highly precise measurement of a source’s DM, which would be indicated by an improved agreement with fits to the sub-burst slope law. We did indeed observe slight improvements in the goodness of fit compared to using each burst’s individual DM. In Sections 3.2 and 4.1.3, we detail the process of applying each burst’s individual DM including, recalculating DMs when necessary. The results of these DM investigations are discussed further in Appendix C.
@@ 3.2 Burst DM Measurements
As mentioned, to assess any major differences in our results due to the choice of DM, we repeat our measurements at the individually determined DM of each burst. We repeated all spectro-temporal measurements at the burst DM reported by the original authors, if there was one. For bursts without an individually calculated DM, we used DM_phase555https://github.com/danielemichilli/DM_phase (CHIME/FRB Collaboration et al., 2019; Seymour et al., 2019) to recalculate the DM using an algorithm that preserves structure by maximizing coherent power instead of S/N, and repeated the measurement at that DM. The datasets whose burst DMs we recomputed were those from Gajjar et al. (2018); Michilli et al. (2018); Oostrum et al. (2020); Scholz et al. (2016); Snelders et al. (2023); Hewitt et al. (2022); Zhang et al. (2023); Nimmo et al. (2022, 2023); Pleunis et al. (2021b); Marthi et al. (2020); Bethapudi et al. (2023). For most of the FRBs we recomputed DMs for, we searched for optimal DMs in a range of 40 pc/cm3 from the applied DM in steps of 0.1 pc/cm3. For shorter duration FRBs (namely those from Snelders et al. 2023 and Nimmo et al. 2022, 2023, as well as bursts from FRB 20180916B), a finer grid was needed over a smaller range, so we searched DMs 10 pc/cm3 from the applied DM in steps of 0.01 pc/cm3. For the remaining datasets that reported individual burst DMs, we repeated our measurements at those DMs.
@@ 4.1 Spectro-temporal relationships
Table 2 presents the results of these fits for each FRB source analysed. The fit for FRB 20121102A, which includes 149 bursts measured and covers the widest range of frequencies and durations, yields and ms. The uncertainties here and throughout (including figures) are quoted at the 1 level. This fit is shown as the black dash-dot line in Figure 4. Note that the curved appearance of the fit line results from the logarithmic scale used on the duration axis. We check the difference in fit when using the average channel duration instead of the integrated time-series duration. With that duration, the fit changes to and ms. We note however that the residuals are larger in this case and tend to larger positive values for longer duration bursts. Due to the general form of the fit, agreement with Equation 4 is qualitatively captured by the magnitude of (as well as ), where a value near zero indicates good agreement and larger values suggest otherwise. Assuming that must be zero (i.e., that the sub-burst relation holds true) allows to serve as an indicator of data quality. A large may result from a limited variety or incomplete sample of bursts in terms of frequency and/or duration. This trend is evident in Table 2 for the remaining sources, where fluctuates around the value found for FRB 20121102A, while generally increases significantly as the number of available bursts decreases. FRB 20220912A challenges this trend since is large despite the large number of bursts, however this may be due to an actual deviation from the sub-burst slope law for this source.
@@ 4.1.2 Relations between , , and
The – plot appears to exhibit a linear correlation, as was found with previous results obtained over broad frequency ranges but with fewer bursts (Houde et al., 2019; Chamma et al., 2023; Brown et al., 2024); however, there again remains significant statistical scatter. Additionally, the higher frequency bursts shown are observed with a larger observing bandwidth. The observing bandwidth may therefore be biasing this result. We fit a line to the complete set of measurements and obtain . Previous values for this correlation are around , after inverting and scaling from the FWHM for comparison here. Despite the present value being nearly double, we note that, because of the spread in measurements, either fit could be visually attributed to the data.
Two fits are overlaid on the – plot of the form . The blue and orange lines are fit to all bursts with durations longer and shorter than 300 s, respectively. We find [msMHz] for the longer FRBs, and [msMHz] for the ultra-FRBs. Despite the statistical scatter in measurements, the blue line reflects well the trend of measurements, especially the very low frequency bursts from FRB 20180916B. The fit to the ultra-FRBs (orange) also describes the shortest duration bursts well. Bursts at around 1.4 GHz, especially from the sources FRB 20200120E and FRB 20220912A, connect the duration gap between the two fits, suggesting a continuum may exist. In that case, the blue and orange fits may represent opposite ends of an envelope. The functional dependence is the prediction of the TRDM and consistent with results reported in Chamma et al. (2023). There is room for other interpretations and functional forms that may describe this data (such as ). However, we focus here on the context that the TRDM provides and follow predictions made by Kumar et al. (2024), which investigates modifications to the TRDM at short durations and under various propagation effects. In Kumar et al. (2024), bursts that obey the sub-burst slope law and vary in duration by several orders of magnitude can show the sort of separation reflected in the two fits shown.
@@ 4.1.3 Changes when measuring at individual burst DMs
Generally, we observe an expected shift of slope measurements upwards towards positive slopes. After applying the same measurement filters applied to the earlier measurements (Section 2.2), 325 sub-burst measurements remain (shown in Figures 11 and 12). The reduction in burst count is primarily due to the increase in uncertainty for values of d/d that are close to or overlap with 0 ms/MHz. Since the burst DM is found using an algorithm that generally aligns bursts vertically, this situation is more common and we observe a corresponding decrease in usable measurements. Nonetheless, we still observe a downwards trend with increasing duration consistent with the sub-burst slope relation. The fit parameters obtained for the sub-burst slope relation for bursts from FRB 20121102A was and ms, as well as a larger reduced-. For the remaining sources, the fit parameters found showed similar differences, with the reduced- either similar or much larger compared to those tabulated in Table 2.
@@ 4.2 Drift rates
Figure 6 shows the drift rate measurements, each multiplied by the mean frequency of their underlying components, plotted against their duration. Overlaid is the fit obtained for the analogous plot of the sub-burst slopes in Figure 4. We observe good agreement between this fit and the drift rate trend, supporting the findings in Jahns et al. (2023) and Brown et al. (2024) that these two measurements follow similar trends. Interestingly, the positive drifts approximately follow the negative of this relationship, shown with the faded line, though the small number of points makes it difficult to conclude for certain. This suggests a connection with the negative drifts and the existence of a sign change (Rajabi et al., 2020), which can explain this aspect of the positive drifts. To double check this result, we repeated all our drift rate measurements using the ACF method. We obtained 132 drift rates from our sample in this way (shown in Figure 13). Using the same durations, we saw similar trends, despite some differences in measured drifts. With the ACF drift rates, the values move above the line whereas they were below in Figure 6. The positive drift rates also exhibit a larger spread in measurements. We note that truncated drift rates are also potentially inaccurate using the ACF method (just as they can be with the arrival times method). Truncated signals can lead to a rotation of the ACF, changing the measured drift rate. Additionally, as the separation between components becomes greater, multiple distinct features can arise in the ACF. This makes the ACF drift rate more challenging and ambiguous to determine as the model may not fit these features well. In this latter case, the arrival times method offers advantages over the ACF method for drift rates.
@@ 5 Discussion
The slope–frequency relation is a strong correlation that is observed in our results here and predicted in the TRDM (Eq. 7 of Chamma et al., 2023). Despite three significant outliers (out of seven total points), the addition of low frequency LOFAR data for FRB 20180916B seems to suggest a related trend between at least that source and FRB 20121102A. This is because the same scaling of the fit shown in the – panel of Figure 5 describes most of the FRB 20180916B data and indeed the data from all sources well. In the TRDM the scaling constant for this relationship is equivalent to where is the rest frame frequency of emission of the FRB. This relationship is also expected in the context of curvature radiation by charged bunches in the magnetosphere of magnetars (Wang et al., 2022). In this vein, a recent scintillation analysis for FRB 20221022A argues that the lateral size of the FRB emission region in that source to be less than km (Nimmo et al., 2025).
@@
We also note that the sub-burst slope of ultra-FRBs is especially sensitive to dispersion errors. Because of their extremely short duration, ultra-FRBs appear close to vertical on a perfectly dedispersed waterfall. It follows that even small levels of over-dedispersion can change the sign of their sub-burst slope (Kumar et al., 2024, their fig. 10). In principle, this could explain the tendency for some of the sub-burst slopes of shorter duration bursts, especially those from FRB 20200120E (durations between and ms) and FRB 20121102A and FRB 20220912A (durations above ms), to ‘bump’ upwards in Figure 4.
@@ 6 Summary and Conclusions
We have developed and applied a measurement pipeline that uses the arrival times of an FRB pulse in each frequency channel of its dynamic spectrum to perform sub-burst slope/intra-burst drift measurements. The pipeline successfully delivered high precision measurements of the slope including for ultra-FRBs. In most cases, earlier 2D Gaussian techniques yielded measurements with similar values and uncertainties, whereas the ACF method yielded the smallest uncertainties across all methods used. Drift rates were also measured for bursts with multiple resolved components, yielding a total of 143 measurements.
@@ Acknowledgements
The authors would like to thank the anonymous referees, whose feedback greatly improved the quality of the manuscript and its conclusions.
@@ Appendix C DM optimization using the sub-burst slope relation
We apply this process for the sources FRB 20121102A and FRB 20220912A separately. Note that we exclude the many bursts from Hewitt et al. (2022) for FRB 20121102A due to their different applied starting DM and burst B1 from Hewitt et al. (2023) for FRB20220912A due to the presence of significant blending. The starting DM used is the DM of the shortest (or nearly the shortest) duration burst from that source. In the case of FRB 20121102A that burst is the microburst B30 of Snelders et al. (2023), and burst B1 of Hewitt et al. (2023) for FRB 20220912A. The results of this are shown in Table 4.
@@
Appendix D Figures with Adjusted Methodologies
We include here Figures 11, 12, and 13. These figures are copies of Figures 4, 5, and 6 but with adjustments in the methodology used. Figures 11 and 12 display measurements from each burst’s individually determined DM, and are discussed in detail in Section 4.1.3. Figure 13 presents multi-component drift rates measured using the ACF method to contrast with the arrival times drift rate measurements. This figure is discussed in detail in Section 4.2.